### v2

Abstract
Digital health tools have the potential to significantly improve the delivery of healthcare services. However, their adoption remains comparatively limited due, in part, to challenges surrounding usability and trust. Large Language Models (LLMs) have emerged as general-purpose models with the ability to process complex information and produce human-quality text, presenting a wealth of potential applications in healthcare. Directly applying LLMs in clinical settings is not straightforward, however, with LLMs susceptible to providing inconsistent or nonsensical answers. We describe how LLM-based systems can utilize external tools and provide a novel interface between clinicians and digital technologies. This enhances the utility and practical impact of digital healthcare tools and AI models while addressing current issues with using LLMs in clinical settings such as hallucinations. We illustrate LLM-based interfaces with examples from cardiovascular disease and diabetes risk prediction, highlighting the benefit compared to traditional interfaces for digital tools.
Digital healthcare technologies represent a frontier in medicine and a multitude of tools have been developed sutton2020overview ; dunn2018wearables , in particular risk scores for prognostic modeling moons2009prognosis . Clinical adoption of such methods, however, faces significant hurdles eichler2007barriers ; Mathews2019 , with some even calling their use “infeasible” Muller2010 and “substantially conceptual” abernethy2022promise . One key issue is usability Ratwani2019decade , which can result in errors associated with patient harm Howe2018electronic and contribute to clinician frustration, jeopardizing patient safety Shanafelt2016 ; gardner2019physician .
New tools employing artificial intelligence (AI) and machine learning offer substantial promise Bajwae2021 . Yet these approaches face the same usability challenges as existing digital tools, while AI systems present additional questions around model trust Rajpurkar2022 ; asan2020artificial . These issues have limited the uptake and impact of AI technologies thus far goldfarb2022ai ; Davenport2019 .
Large Language Models (LLMs) offer a potential solution to the challenges faced by digital tools and risk scores if used appropriately. LLMs have recently captured the imagination of both the research community and the general public, pushing the boundaries of human-machine interaction. Consequently, there is great interest in applying LLMs in healthcare, with potential applications including facilitating clinical documentation, summarizing research papers, or as a chatbot for patients Moor2023 .
Figure 1: Clinicians have previously needed to interact directly with digital tools, such as risk scores. While others have discussed LLMs replacing existing clinical tools, we envisage LLMs forming a novel interface with clinical tools by providing a dynamic output and facilitating deeper interaction with the tools and related information, such as model explainability, medical papers, and clinical guidelines.
Applying LLMs in safety-critical clinical settings is not straightforward. LLMs may provide inconsistent or nonsensical answers Singhal2023 ; Lecler2023revolutionizing and have a tendency to hallucinate facts maynez2020faithfulness ; Ji2023survey . This is clearly unacceptable in medicine when making high-stakes decisions. Additionally, LLMs can encounter difficulty with seemingly basic functionality, such as mathematical calculations or factual lookup patel2021nlp ; schick2023toolformer , and are unable to access up-to-date information by default komeili2022internet . These limitations constrain the clinical utility of LLMs.
In this Comment, we explore a new application of LLMs in healthcare and propose using LLMs to facilitate clinician interactions with AI models and digital tools. Conceptually, this differs substantially from previous applications, such as training medicine-specific LLMs Luo2022 ; Yang2022large or using LLMs for prediction Jiang2023health . Instead, we construct LLM-based systems, offering an intuitive natural language interface that can streamline clinician interactions with multiple digital tools and sources of information, improving efficiency and usability (Fig. 1). By adopting a systems approach, since the LLM does not itself issue predictions or supply specific knowledge, the potential for hallucinations is limited and we ensure actionable information is provided by approved clinical sources.
LLM-based user interfaces can help engender trust in digital tools and AI models, such as risk scores, by enabling dynamic interactions in the form of natural language dialogues, which has been identified as a key feature for practitioners to deploy machine learning models in healthcare lakkaraju2022rethinking . The most appropriate interface or dialogue cannot be pre-specified; instead, it depends on the clinician and the patient. Thus, the ability to adapt and tailor interactions represents a critical advance in the functionality of such tools that can be unlocked by LLMs.
Table 1: Representative questions that a clinician might have relating to a risk score at different stages of a patient encounter, together with whether existing interfaces for risk scores provide this information. All questions can be addressed using LLM-based interfaces.
Existing
Stage Representative questions interfaces
Before Patient
Which features does the risk score use?
Encounter
Why are these features included in the risk score?
How was the risk score validated?
What is the methodology underlying the risk score?
Before Risk
When do clinical guidelines recommend risk scoring?
Scoring
What is the recommended risk score?
Who is the risk score suitable for?
During Risk
What is the risk for this patient?
Scoring
What characteristics led to the patient’s risk?
What effect would changing this feature have on the risk?
After Risk Scoring
What action is recommended by the guidelines based on the risk?
As a concrete example, let us consider risk scores. Currently, clinicians access these tools via fixed user interfaces or application programming interfaces (APIs). Numerous considerations exist before, during, and after calculating an individual’s prognosis (Table 1), while existing interfaces only serve to calculate risk. Through an LLM-based interface, practitioners can obtain substantial additional information about the risk score, its development and methodology, the prediction issued, and related medical guidelines, in a manner that specifically addresses their needs or questions without providing superfluous information.
By default, LLMs do not possess the ability to access external tools or information. Methods to extend the functionality of LLMs beyond text generation are in their infancy but already enable powerful LLM-based systems. Below, we describe such a system, illustrated in Fig. 2. Instead of solely relying on an LLM to issue predictions or provide information, we can augment LLMs and enable them to access approved medical tools and other sources of information, thereby not solely relying on the inherent capabilities of a given LLM. This approach is scalable to multiple predictive models, unifying digital tools within a single, natural language-based interface. This combines the strengths of LLMs with digital healthcare tools, while overcoming limitations of both, and contrasts previous applications of LLMs in medicine.
1 Illustrative LLM Interfaces
To illustrate the ideas described in this paper, we built example instantiations of LLM-based interfaces for cardiovascular disease (CVD) and diabetes risk prediction (Fig. 3). To demonstrate the versatility of our approach, we took an off-the-shelf pretrained LLM, GPT-4 openai2023gpt4 , and imbued it with a set of tools using in-context learning, requiring no additional training of the LLM. Additionally, we implemented a “source” functionality that allows the user to see whether the LLM used a tool or accessed specific information and, if so, which tool and with what input. This helps avoid hallucinations since it enables verification that the information was issued by an underlying clinical tool or source rather than being generated by the LLM. Additional implementation details are provided in Appendix A.
CVD Risk Prediction. We first show how LLMs can incorporate existing tools and information for CVD risk prediction (Fig. 2(a)). We provided the LLM access to QRisk3 Hippisley-Cox2017CVD , a risk prediction tool that assesses the likelihood of developing CVD within 10 years. Additionally, we provided the LLM with access to the academic paper describing QRisk3 Hippisley-Cox2017CVD and the National Institute for Health and Care Excellence (NICE) clinical guidelines for CVD NICE_CVD . An example interaction between a physician and the LLM-based system is shown in Fig. 2(a). An illustration of the reasoning process by which the LLM uses external tools is provided in Fig. S.1).
Figure 2: Overview of an LLM-based system that enables clinicians to interface with digital tools using natural language inputs. (1) The LLM is provided with the history of the interaction, including the current request. (2) Using an iterative reasoning process, the LLM decides which, if any, tools are required and with what input. (3) The LLM provides a response to the user incorporating information provided by any tools that were used.
In this example, when asked for the patient’s 10-year risk of developing CVD, the LLM uses QRisk3 to estimate the patient’s risk, providing this to the user. The LLM then summarizes the QRisk3 paper to explain the inclusion of certain features before providing the recommended action for this patient from the NICE clinical guidelines. Finally, the LLM uses QRisk3 to recalculate the patient’s counterfactual risk assuming that they were able to reduce their systolic blood pressure to within normal ranges. This allows both the clinician and patient to understand the potential impact of changes to modifiable variables on the patient’s risk. While the clinician could have used the underlying resources to obtain this information, the LLM interface made the interaction simpler and more efficient, which has been identified as a key limitation of digital tools Ratwani2019decade ; Mathews2019 . In the following example, we will explore a more dynamic output extending beyond model use and information retrieval to new functionality.
(a) Cardiovascular disease risk prediction
(b) Diabetes risk prediction
Figure 3: Example interactions with LLM-based interfaces. (a) CVD: The LLM uses QRisk3 to calculate the patient’s 10-year risk of developing CVD, explains why certain features are included in the risk score using the QRisk3 paper Hippisley-Cox2017CVD , and provides recommendations for this patient from the NICE clinical guidelines. (b) Diabetes: After calculating the patient’s risk using the AutoPrognosis model, the clinician can query why this prediction was issued using explainable AI and understand the underlying methodology.
Diabetes Risk Prediction. We now demonstrate an application for diabetes risk prediction using a machine learning model (Fig. 2(b)). We trained a prognostic model using AutoPrognosis Imrie2023AutoPrognosis , an open-source automated machine learning framework, on a cohort of individuals from UK Biobank Sudlow2015 to estimate a patient’s 10-year risk of developing diabetes. Building model trust is a crucial step for prognostic models, in particular for models that are not inherently interpretable Rajpurkar2022 . A recent study found that medical decision-makers had a strong preference for interactive explanations and, in particular, for these interactions to take the form of natural language dialogues lakkaraju2022rethinking . To enable this, in addition to allowing the LLM access to the model, we provided the LLM with the ability to apply explainable AI (XAI) methods Imrie2023 on the underlying model, allowing users to investigate the rationale for predictions.
As shown in Fig. 2(b), after calculating the patient’s risk using the AutoPrognosis model, the clinician can query why this prediction was issued using XAI, facilitating a deeper understanding of the model. In this example, we demonstrate the use of Shapley additive explanations (SHAP) lundberg2017unified . Due to their general pretraining, LLMs have knowledge of many topics; this can provide valuable additional information during interactions beyond the specific tools and external information sources provided to the LLM. For example, suppose the clinician is not familiar with the underlying XAI methodology, SHAP lundberg2017unified . The LLM is able to explain how this approach works, in a variety of different ways and possibly over multiple interactions with the clinician, allowing specific queries or misunderstandings to be clarified. This demonstrates the utility of LLMs beyond simply using existing tools and illustrates the benefit of their underlying knowledge.
2 Discussion
Large Language Models hold substantial promise for the medical domain, particularly in augmenting digital workflows and improving efficiency in healthcare delivery. The ability to integrate external tools and functionality with LLMs paves the way for innovative applications and can overcome limitations of LLMs, such as the potential to hallucinate.
We have demonstrated how LLMs can provide a unique interface between healthcare professionals and clinical predictive models, such as risk scores. Our approach is scalable and requires no additional training, although approaches that improve with use could be yet more powerful. Additionally, we aim to mitigate the problem of LLM hallucination and ensure that actionable advice is anchored in approved clinical resources.
While LLMs have general capabilities, they can lack domain-specific knowledge. This has led to the development of medical-focused LLMs, either by training new LLMs from scratch Luo2022 ; taylor2022galactica ; Yang2022large or by adapting existing general-purpose LLMs Singhal2023 . While we showed using such specialist LLMs is not required, they are easy to incorporate due to the modularity of the approach and doing so could further enhance the functionality of LLM-based interfaces.
As AI in medicine continues to advance, further research into LLMs and their potential applications is not only beneficial but vital. LLMs could help to alleviate the data burden that is contributing to clinician burnout, as well as streamline patient management processes. Furthermore, studies have demonstrated high usability of LLMs, even with limited experience Skjuve2023user , which is critical for successful clinical deployment. While we believe this paper represents an important first step, we are only scratching the surface of the potential of LLMs in healthcare. Ultimately, this line of work may significantly change the digital health landscape, enhancing the capabilities of clinicians and the quality of patient care.
3 Code availability
Code for the LLM-based interfaces can be accessed at https://github.com/pauliusrauba/LLMs_interface.
4 Data availability
The risk predictor for diabetes was trained on a cohort from the UK Biobank resource. Data from UK Biobank is accessible through a request process (https://www.ukbiobank.ac.uk/enable-your-research/register). The authors had no special access or privileges when accessing the data.
[…]
Acknowledgments This study received no funding.
[…]
Competing interests All authors declare no financial or non-financial competing interests.
[…]
Author contributions F.I. and M.vdS. conceptualized the manuscript. F.I. and P.R. designed and performed the experiments. F.I. wrote the original draft of the manuscript, and all authors contributed to editing and revising it.
    A.1 Large Language Models
    By default, LLMs will provide responses in the form of text based on the provided context, such as a prompt or conversation history. To construct interfaces for digital tools using LLMs, we instead view the LLM as an agent that can interact with an environment to solve tasks. Formally, at each step $t\in T$, the agent receives observation $o_{t}\in O$ from the environment and subsequently takes action $a_{t}\in A$ according to policy $\pi(a_{t}|h_{t})$, where $h_{t}=(o_{0},a_{0},\dots,o_{t-1},a_{t-1},o_{t})$ is the history. To enable the agent to both reason and use external tools, we use the ReAct framework yao2022react which decomposes the action space as $\hat{A}=A\cup L$, where $a\in A$ are actions using specific tools and an action $a\in L$ represents not using an external tool but instead allows the model to reason over the history about what action to take next.
    Since we will provide the agents tasks in the form of natural language, and actions in the language space $L$ are (essentially) infinite, we chose to benefit from strong language priors and use a pretrained LLM. We implemented our LLM-based interfaces using GPT-4 openai2023gpt4 , with access through the OpenAI API. Interactions with external tools were implemented using LangChain LangChain .
    While frameworks such as Toolformer schick2023toolformer and WebGPT nakano2021webgpt trained LLMs to use basic tools, such as calculators, calendars, and search engines, via self-supervised fine-tuning and fine-tuning using behavior cloning and reinforcement learning, respectively. Following ReAct yao2022react , we used in-context learning dong2023survey in the form of prompting, providing the LLM with sufficient information about possible actions and using the underlying reasoning capabilities of the LLM. This removes the need for further training of the LLM, which might be challenging in the medical domain without suitable examples. This also enabled us to use multiple tools. We built user interfaces using StreamLit streamlit .
    A.2 External Tools
    This section describes the external tools and sources of information made available to the LLM in our two examples. In the subsequent section, we provide the system and tool prompts.
    CVD Risk Prediction. In our first example, we show how LLMs can incorporate existing tools and information for CVD risk prediction (Fig. 3a). We provided the LLM access to QRisk3 Hippisley-Cox2017CVD , a risk prediction tool that assesses the likelihood of developing CVD within 10 years. QRISK3 is currently the recommended risk score for CVD in the United Kingdom and is widely used to guide decisions about who might benefit from interventions NICE_CVD . We enabled the LLM to use the risk score either using the provided data or, if requested by the user, to modify a variable and assess the impact of such a change on the patient’s risk. Additionally, we provided the LLM with access to the academic paper describing QRisk3 Hippisley-Cox2017CVD and the National Institute for Health and Care Excellence (NICE) clinical guidelines for CVD NICE_CVD .
    Diabetes Risk Prediction. To demonstrate an application for diabetes risk prediction using a machine learning model (Fig. 3b), we equipped the LLM with the following tools. For the diabetes risk score, we trained a predictive model using AutoPrognosis 2.0 Imrie2023AutoPrognosis on a cohort of individuals from UK Biobank Sudlow2015 with no known history of diabetes at baseline. The primary outcome was diagnosis of diabetes within a 10-year horizon. We selected 20 features based on their predictive value Imrie2023AutoPrognosis , namely HbA1c, blood glucose, sex, age, waist size, hip size, weight, BMI, waist-height ratio, waist-hip ratio, gamma-glutamyl transferase, cystatin-c, C-reactive protein, atrial fibrillation, HDL, alanine transaminase, triglycerides, uric acid, sex hormone-binding globulin, use of anti-hypertensives.
    Building model trust is a crucial step for prognostic models, in particular for models that are not inherently interpretable Rajpurkar2022 ; asan2020artificial . A recent study found that medical decision-makers had a strong preference for interactive explanations and, in particular, for these interactions to take the form of natural language dialogues lakkaraju2022rethinking . To enable this, we enabled the LLM to use explainable AI (XAI) methods Imrie2023 on the underlying model, allowing users to investigate the rationale for predictions, both in general and for the specific patient. In particular, we used SHAP lundberg2017unified to interpret model predictions.
    Figure S.1: Illustration of the process by which the LLM uses external tools. The LLM is provided with the history of the interaction, including the current request. Using an iterative reasoning process, the LLM decides which, if any, tools are required and with what input (“Thought”). The LLM then uses the external tool (“Action”) and receives the output (“Observation”). Finally, the LLM decides to answer the question (“Thought”), providing a response to the user.